name2004-iss4: Logbook Entries

name2004-iss4: INFO Messages: 9 Entries..

Return to Logbook Contents Page
Entry Date Title Site Author #Graphics
254 Wed 04-Aug-2004Los Mochis Roving crew visitISS4 (Los Mochis)June
244 Tue 03-Aug-2004Roving crew visit: problems with losing data earlierISS4 (Los Mochis)June
243 Tue 03-Aug-2004LORETO Roving crew visitLou
190 Tue 27-Jul-2004Problems for 27 July 04 Morning CrewISS4 (Los Mochis)Nick
116 Sat 17-Jul-2004Roving team notes for July 17ISS4 (Los Mochis)Steve/John
108 Fri 16-Jul-2004Roving team notes for July 16ISS4 (Los Mochis)Steve/John
99 Thu 15-Jul-2004Roving team notes for July 15ISS4 (Los Mochis)Steve/John
98 Thu 15-Jul-2004DATA BACKUPS: PLEASE READISS4 (Los Mochis)Steve/John
80 Wed 14-Jul-2004Roving crew visitISS4 (Los Mochis)Steve/John

254: INFO, Site ISS4 (Los Mochis), Wed 04-Aug-2004 18:05:52 GMT, Los Mochis Roving crew visit
The roving crew (Lou and June) visited Los Mochis-ISS4 site from August 2 to 
August 5. Here are the summary of our visit:

(1) Investigation of "408" sondes: After looking at all sondes failed to launch 
at both Los Mochis and Kino, they all are 4084141XX (manufactured on 2004, week
 08, day 4 on calibration tray 141). All operators were informed not using those
 sondes. We have total 12 4084141XX sondes: six attempted to launch at LM, two 
at Kino, three at LM not launched, and one unaccountable (408414100). All 12 
bad sondes at LM were packed in a box labeled with "Prob. sondes". We compared 
the coefficients of one sonde (408414113) from the paper slip with that on 
disk. They are the same. 

(2) Backup data on DM computer: We backed up data on DM computer on two DVDs 
(one for raw data and one for Nima), left one copy in the trailer and 
brought one copy with us.

(3) Backup profiler data: We stopped the profiler at 10:40AM-12AM LST to backup
 the data. We backed up data to two CDs (NAME_LMM03 and NAME_LMM04), left one
 copy in the trailer and brought one copy with us. After backup, the disk space
 is ~8.5% full.

244: INFO, Site ISS4 (Los Mochis), Tue 03-Aug-2004 18:31:27 GMT, Roving crew visit: problems with losing data earlier
After looking at some of soundings losing data earlier, I think that there are 
two general reasons for losing data earlier:

(1)Telemetry problem: If there are telemetry problems, you lose PTU data 
earlier, but still have wind data, so data wouldn't show up in plots without 
pressure data. The sounding at 08Z on July 29 at Los Mochis show no PTU data in 
700-180 mb, but have wind data.

(2)Possible problem with post-processing software: if losing data for a period 
of time, the data are still transfering, and everything displays properly. The 
sondes were broken above 100 mb. After you run the postprocessing, you only see
 data below ~400 mb. We should investigate this more after looking at all data.


243: INFO, Site , Tue 03-Aug-2004 17:13:12 GMT, LORETO Roving crew visit
July 31st and August 1st the roving crew spent at the Loreto site. The first job was to look at the Motosat 4020 system there. The problem appears to have started when by mistake the motosat admin software was closed by accident. This should not be a problem as the software can be turned off with the the dish still deployed.The problems started when the software was restarted at which time the
dish stowed this should not have happened.Also the Motosat controller could not be shut off.Open this box to see if the switch was not being activated.Switch appears fine when turned off data and dish lights go off for approx 1 - 2 minutes
then come back on,LNB and blue power switch remain on. This could be affecting the disk as to emergency stow the dish the power switch can be pushed until dish is stowed.Also tried to test the dish by appling dcv to the elevation lines leading into the controller box normally 9 volts is applied to move the dish but at Franks suggestion[Eagle Computers] also tried a 12vdc source since this is 1.2 
meter dish.Could not get any movement from dish or even sounds of the motor tring to move. Removed cable took it in and checked all lines. No opens or shorts on any of the leads. In discussions with Frank came to the conclusion there maybe
multiple problems one being dish moter the other being controller box. [NB we 
had to replace one controller box before this on the MISS system supposedly there was a bad run of controllers according to Eagle. 
On a better note there have been no appearant problems with the GAOS and 2 ISS 
4000 series systems at the other sites.
Since the Motosat will not be able to be repaired for the porject the students will attemp once a day to down load data via an Internet cafe [these appaer to be closed on sundays]. This could create a possible problem for late notification of IOPs or extentions of IOP's.
Secondly at LORETO we replaced the questionable sounding computer with one
brought down for that reason after June cleaned up the hard drive she also ran the last 5 days worth of data through ASPEN to create the temp files. June backed up and looked at all the sounding data
We also tried to move the sounding system telemetry antenna. A 100foot cable was used and during a flight was moved to several locations. The original location gave far higher, 25-30 db reading on receiver.

190: INFO, Site ISS4 (Los Mochis), Tue 27-Jul-2004 11:36:50 GMT, Problems for 27 July 04 Morning Crew
During the 1200z launch we encountered a problem.  We opened a new box of sondes that began with the coefficient of 408.  I used these at the Kino Bay site with no problems.  We were able to pass through the frequency scan and the coefficients were accepted.  The problem is that we were receiving readings on the order of 1080mb of pressure and 80C temp readings.  We were also picking up a 99db signal prior to hooking up the battery at the 403Mhz frequency.  There was also a warning tag at the bottom of the window where the met data begins to display that stated, "WARNING! This is RAW DATA."  Neither myself or Nick have seen this and we speculated that the error was coming from the sonde itself.  We attempted to use 2 sondes with the 408 coeffs and decided to try a third with a different coeff since they have not used the 408 coeffs here up to this point.  We used the same battery for all three sondes so we still have the capabilities of using the inital two sondes if we resolve this problem. Please let us know of any possible solutions and we will use the two initial sondes first.  It is a concern to us since we are now down to 98 sondes at this location. 


116: INFO, Site ISS4 (Los Mochis), Sat 17-Jul-2004 14:34:14 GMT, Roving team notes for July 17
We got through a few more tasks today.

Backup and deletion of ALL profiler days through yesterday. Also cleaned
up the D: drive. There is now space for POP to run for 23 days (through
Aug 7) but we should backup more often so well before then.
Profiler was down from 1405-1435Z for this backup.

We did at test, swapping 6L/6H and back. Results: the IR dropped a little
from say 440 to 415, but it is overcast and not a good test. Changed back
and we will repeat this test at another site, another day.

GLASS preamp:
Power draw from Campbell battery to CLASS preamp-box is continuous for now.
Over night the Campbell battery voltage dropped slightly to 12.85V and
recovered to 13.05V. It is not recovering fully to its values from past days.
We will monitor this. If needed, this power link can be switched off when
not in use. If this task turns out to be too much for the Campbell, we will
have to come back and wire in the 12V external power supply.

PC-208W: Copied software onto a CD in case we need it at in Loreto.

108: INFO, Site ISS4 (Los Mochis), Fri 16-Jul-2004 22:15:41 GMT, Roving team notes for July 16
We got through a few more tasks today.

Using the Campbell to power the CLASS Rx pre-amp worked well on the past 4
soundings. Also, there was only a small drop in Campbell battery voltage
over the past 24 hours. JWM neatened the new wiring connection. There is now
an on/off switch in the Campbell box. A small hole was drilled in the CLASS
Rx box. We asked Gary to add V_batt to a web plot so we can monitor this

We made 2 copies of SPC files for Days 179-193. Verified them as readable
(at least a few files) on laptop. Then deleted these days from the profiler
disk. This disk was 57% full, after deleting was 33% full.

Profiler was down 21:35 to 22:25 for this backup.
Coarse estimate:
  Current disk free space is 1.32GB. SPC files at 73 MB/day. In a pinch, the
disk could hold data for 2.5 more weeks (but better to back up more often).

SURFACE IR SENSOR: Good progress today. A small change in a DLD coefficient
made a large difference in the data. It may be "fixed" but a question remains
about the Campbell wiring. See SURFACE tklog entry#110.

99: INFO, Site ISS4 (Los Mochis), Thu 15-Jul-2004 21:00:28 GMT, Roving team notes for July 15
ISS-4 system and data backup done today. Did not get to profiler PC data backup.

Based on previous notes/theory, the IRT in a cloud free sky should see a
temperature of about -50 to -30 C which becomes a reading of 140-200 W/m2.
However, that may not be the appropriate sky temperature for the current
environment. For OO and Kino "typical" clear sky values are around 350-450 W/m2
indicating a sky temperature of 290-273= +17 C (for 400 W/m2).

We have experimented, covering the sensor with a cardboard box. If the box were
a blackbody we can estimate the IR. When first covered, the IR became about
100 W/m2, corresponding to T=-70 C. This soonincreased to 200 W/m2
corresponding to -30 C. I would expect much higher values if we were seeing the
cardboard box temperature (on the order of +30 C).
Another experiment was removing the battery. That is something Hal had replaced.
No effect. We reinstalled the battery at 22Z today.
Will pursue this more tomorrow.

See JWM excellent summary tklog entry#89. We made a change and will monitor
results. Note that Campbell Tbox must be monitored also. More tomorrow.

98: INFO, Site ISS4 (Los Mochis), Thu 15-Jul-2004 19:53:20 GMT, DATA BACKUPS: PLEASE READ
Please append all backup DVD notes to this log entry.

Performed ISS-4 system backup (no data) on 15-July-2004 (around 19:30Z)

Performed ISS-4 data backup on 15-July-2004 (around 20:45Z)

80: INFO, Site ISS4 (Los Mochis), Wed 14-Jul-2004 02:29:18 GMT, Roving crew visit
John and Steve arrived at Los MOchis this evening. A few initial checks with
more tomorrow.

Profiler level looks good.
NetRad leveled. Was off by 1/2 bubble
PYG from Kino installed. Pins not yet checked, but data is on as of 02:30Z July 14

NOTE: PYG was removed July 8. From then to now the circuit was open and "numbers" appear but they are random and false data.